# Distilling Large Language Models for Network Active Queue Management

### v1

@@ III-B2 Evaluation of Accuracy Dynamics in L4S-LLM
The evolution of the accuracy of L4S-LLM with four different LLM models - Llama2 (7b), OPT (1.3b), GPT2 (125m) and T5-LLM (220m), during the fine-tuning phase is depicted in Figure 8. The Llama2’s accuracy starts near zero, improves significantly between epochs 5 and 9, then plateaus before a gradual increase from epoch 30, gradually stabilising around 97.56%. OPT starts with higher accuracy than GPT2 and Llama2, increasing till epoch 40 where it stabilises to around 83% accuracy. GPT2’s accuracy remains near zero till epoch 29 before it sharply increases till epoch 40 before it stabilizes. T5-LLM starts from 20% accuracy and sees a gradual increase and only after epoch 75 does it seem to stop increasing and stop around 83% accuracy similar to the final accuracy scores of OPT and GPT2.
The results show that Llama2 demonstrates the best performance considering its high training time, compared to the others. OPT, GPT2 and T5-LLM all reach the same accuracy of 82% at the end of epoch 80. It demonstrates Llama2’s superior performance, considering its 7B parameter size, which when combined with LoRA, reduces the model to approximately 70 million parameters for fine-tuning. Despite the high training time of over 20 hrs compared to other models, it exhibits comparable answer generation times in actual deployment to other large language models (LLMs). Thus, not losing any time in actual deployment. Therefore, we use Llama2 as our foundational model for evaluation against real-world scenarios.
@@ III-C Evaluation of L4S-LLM over AQM
In this simulation, we evaluate the performance of our L4S-LLM framework, with Llama2 -7B as our base, fine-tuned using LoRA, with its performance benchmarked against real-world L4S AQM data. The benchmarking focuses primarily on optimizing queue delay and bandwidth utilization. The network topology depicted in Figure 6 involves Client 1 using Cubic and New-Reno concurrently, while Client 2 is running DCTCP, and the router is configured to implement L4S. Using real-world data LLM predicts the AQM action - Enqueue, Dequeue, or ECN marking. The LLM is used at specific intervals as it is resource-intensive. The figures below demonstrates the LLM’s ability to optimize the queue delay and bandwidth according to our reward function.
@@ III-C1 Analysis of Queue Delay
(a) Queue Delay with 10-Interval LLM Inputs
(b) Queue Delay with 100-Interval LLM Inputs
@@
(a) Box Plot of Queue Delay with 10-Interval LLM Inputs
(b) Box Plot of Queue Delay with 100-Interval LLM Inputs
Figure 10: Box Plot of Queue Delay Analysis with Periodic LLM Inputs
Figure 9 presents a comparison of queue delays between the original L4S AQM implementation and the enhanced L4S-LLM framework leveraging the fine-tuned LLM model to optimize queue delay. The LLM model determines the most optimal action – Enqueue, Drop and ECN marking – based on the system’s state. A reward function guides the LLM in finding the most optimal action which keeps improving the reward. The results show a significant reduction in queue delay for the LLM-based system compared to the traditional algorithm. The traditional queue delay has a higher median queue delay with high variability. Conversely, the L4S-LLM achieves a remarkably low median queue delay value, with low variability and fewer outliers, demonstrating the LLM’s capacity to maintain low-latency behaviour. These findings highlight the LLM’s effectiveness, compared to its conventional counterpart, in optimizing dynamic network traffic, ensuring efficient congestion management and enhancing network responsiveness.
@@
Figure 11: CDF of Queue Delay with Periodic LLM Inputs at 10 and 100-Packet Intervals.
@@ III-C2 Analysis of Bandwidth utilization
Figure 12 and Figure 13 showcase the comparative analysis of bandwidth utilization of the traditional L4S and L4S-LLM framework. The results show a significant increase in median bandwidth utilization, with significantly low variability, although along with a few large outliers. We found that these outliers mainly appear at the start of the data transfer and are not present after a certain duration. In contrast, the traditional L4S algorithm in Figure 14 shows a low median bandwidth utilization value with high variability. Similarly, L4S-LLM shows its effectiveness in optimizing bandwidth utilization while maintaining low queue delay. From the observations, we can conclude that LLM’s high consistency and high adaptability in queue management, especially on unseen data.
@@ III-D Computation Overhead(s)
Figure 15 illustrates the number of parameters fine-tuned when using different LLM models. We have utilized LoRA to reduce the number of fine-tuned parameters, as both models originally contain over a billion parameters, significantly impacting training time. Llama2 originally comprised 7 billion parameters, but using LoRA with a rank of 128, the number of trainable parameters is reduced to approximately 70 million, while maintaining performance nearly identical to full fine-tuning. Similarly, OPT, which initially contains 1.3 billion trainable parameters, is reduced to 25 million parameters using LoRA of the same rank. In contrast, GPT-2 and T5-LLM, with 125 million and 220 million parameters, respectively, were used without parameter reduction. The rank value of 128 was chosen because it effectively balances parameter efficiency and model performance, enabling comparable results to full fine-tuning while significantly reducing computational and memory requirements.
@@
Additionally, we also measure total GPU VRAM utilization, which is crucial as it stores the large deep neural network with billions of parameters, numerous layers, and gradients. Insufficient memory limits batch size and slows down both training and inference. Given the high cost of GPU VRAM, research into reducing GPU memory utilization is of paramount importance, as we do not expect routers to be equipped with expensive GPUs containing large amounts of VRAM. Figure 16 illustrates the GPU memory utilization across the respective LLMs. Among the four LLMs used with our L4S-LLM framework, Llama2 occupies the highest GPU memory at 23.91 GB, followed by OPT with 7 GB, GPT2 with 3.6 GB, and T5-LLM with the lowest memory usage at 2.32 GB.
@@
Furthermore, we observe the impact of L4S-LLM on answer generation time as illustrated in Figure 17. The full-parameter, unadapted Llama2-7B has the highest answer generation time of 0.200 seconds, significantly higher than the other methods. The answer generation time for Llama2 is 0.0391s , followed by of OPT at 0.0395s, GPT2 at 0.0414s and T5-LLM at 0.0411s. These answer generation times are largely correlated with the number of trainable parameters and other environment conditions of the system. Though, Llama2 -7B and OPT 1.3B have originally contain over a billion parameters, we are able to reduce it using LoRA which reduces its answer generation time to approximately 0.039s , showcasing substantial performance optimization. This illustrates the efficiency of L4S-LLM in reducing answer generation time while preserving model performance, which is particularly important for latency-sensitive applications. This is a great improvement, but there is more room for improvement. One future direction for improvement is through small language models (SLMs), which have mainly adapted to use in mobile devices. Other solutions include pruning, model partitioning and hardware optimizations.
@@
In Figure 18, we observe the total training time required to run 80 epochs for four different LLMs. Even with the application of LoRA, Llama2-7B exhibits the highest training time by a significant margin, approximately around 22.3 hours. In contrast,the training time for OPT with LoRA is around 29.9 minutes, while GPT2 requires approximately 17.96 minutes, and T5-LLM requires about 32.26 minutes. With the exception of Llama2, the training times for other LLM models are notably faster. However, during evaluation or in actual deployment, the answer generation time is at most importance. Though Llama2’s higher training time, its answer generation performance is comparable to other models.
III-E Discussions and Limitations
Several critical limitations must be addressed before considering the commercial deployment of the proposed L4S-LLM framework in network routers. This section outlines the primary challenges affecting deployability and overall effectiveness:
  • Online Training/Testing: The framework requires real-time evaluation within the L4S architecture to address deployment challenges. While the current L4S-LLM framework was trained and tested using a data-driven approach with experimental data from L4S DualPI2 AQM and various TCP congestion control algorithms, this process involved offline training and testing. Real-time deployment within a virtual router demands further optimization of response generation times and consideration of the resource constraints of the target routers.
@@ III-E Discussions and Limitations
  • Energy: Continuous operation of LLMs requires substantial energy, typically supported by high-performance GPUs and CPUs. The significant energy demands and associated costs pose a challenge to commercial viability. Major technology firms are exploring nuclear-powered energy solutions to meet their LLM needs. Achieving commercial feasibility will require balancing enhanced performance with reduced energy and operational costs, necessitating further advancements.
@@ IV Conclusion
This research develops and presents AQM-LLM, a framework that adapts LLMs for AQM and evaluates it within the L4S architecture. L4S-LLM dynamically manages congestion, optimizes queue performance and minimizes latency through preemptive ECN signalling and packet-dropping mechanisms. The framework achieves significant computational efficiency using low-rank adaptation, reducing memory usage by 64%, and a data-driven reinforcement learning pipeline, which streamlines fine-tuning. The open-source experimental platform of L4S-AQM on FreeBSD-14 facilitates AQM-LLM future research and assists IETF recognition. Although L4S-LLM demonstrates transformative improvements in latency, throughput, and efficiency, further optimization through techniques like pruning, quantization, and distillation toward Small Language Models based AQM is necessary to address resource constraints, paving the way for scalable and intelligent queuing solutions.

### v3

@@ III-B2 Evaluation of Accuracy Dynamics in L4S-LLM
Figure 10 illustrates the training accuracy evolution across four LLM architectures during fine-tuning. Llama2 exhibits initial near-zero accuracy with rapid improvement between epochs 5-9, followed by plateau and gradual ascent from epoch 30, ultimately achieving superior performance at 97.56%. OPT demonstrates higher initial accuracy than competitors, steadily increasing until epoch 40 where it stabilizes at 83%. GPT2 remains stagnant near zero until epoch 29, then exhibits sharp improvement through epoch 40 before stabilization. T5-LLM begins at 20% accuracy with gradual progression, plateauing around 83% after epoch 75.
Llama2 significantly outperforms all alternatives, achieving 97.56% accuracy compared to the 82-83% plateau reached by OPT, GPT2, and T5-LLM. Despite requiring over 20 hours training time versus competitors’ sub-hour requirements, Llama2’s superior accuracy justifies the computational investment. Critically, LoRA reduces Llama2’s 7B parameters to 70M trainable parameters while maintaining performance, and inference times remain comparable to smaller models during deployment.
@@ III-C Evaluation of L4S-LLM over AQM
We evaluate the performance of our L4S-LLM framework, with Llama2 -7B as our base, fine-tuned using LoRA, with its performance benchmarked against real-world L4S AQM data. The benchmarking focuses primarily on optimizing queue delay and bandwidth utilization. The network topology depicted in Figure 8 involves Client 1 using Cubic and New-Reno concurrently, while Client 2 is running DCTCP, and the router is configured to implement L4S. Using real-world data LLM predicts the AQM action - Enqueue, Dequeue, or ECN marking. The LLM is used at specific intervals as it is resource-intensive. Our results, Figs. 11 and 12, demonstrate the LLM’s ability to optimize the queue delay and bandwidth of both DCTCP and UDP Prague flows.
@@ III-C1 Analysis of Queue Delay
(a) Aggregate Delay with 10-Interval LLM Inputs with DCTCP
(b) Aggregate Delay with 100-Interval LLM Inputs with DCTCP
(c) Classic Queue with 10-Interval LLM Inputs with UDP Prague
(d) Classic Queue with 100-Interval LLM Inputs with UDP Prague
@@
(a) DCTCP: 10-Interval Inputs
(b) DCTCP: 100-Interval Inputs
(c) UDP Prague: 10-Interval Inputs
(d) UDP Prague: 100-Interval Inputs
Figure 12: Box Plot of DCTCP and UDP Prague Queue Delays of Fig 11.
Figure 11 presents a comparison of queue delays between the original L4S AQM implementation and the enhanced L4S-LLM framework leveraging the fine-tuned LLM model to optimize queue delay. The LLM model determines the most optimal action – Enqueue, Drop and ECN marking – based on the system’s state. As shown in Figure 12, our reward function guides the LLM in finding the most optimal action which keeps improving the reward. The results show a significant reduction in queue delay for the LLM-based system compared to the traditional algorithm. The traditional queue delay has a higher median queue delay with high variability. Conversely, the L4S-LLM achieves a remarkably low median queue delay value, with low variability and fewer outliers, demonstrating the LLM’s capacity to maintain low-latency behaviour. Our extensive analysis hereafter primarily focuses on DCTCP/VirtualBox, though results for UDP Prague/VMware have been produced using the code available in our GitHub repository.
@@
Figure 13: CDF of DCTCP Queue Delay with Periodic LLM Inputs.
@@ III-C2 Analysis of Bandwidth utilization
Figure 14 and Figure 15 showcase the comparative analysis of bandwidth utilization of the traditional L4S and L4S-LLM framework. The results show a significant increase in median bandwidth utilization, with significantly low variability, although along with a few large outliers. We found that these outliers mainly appear at the start of the data transfer and are not present after a certain duration. In contrast, the traditional L4S algorithm in Figure 16 shows a low median bandwidth utilization value with high variability. Similarly, L4S-LLM shows its effectiveness in optimizing bandwidth utilization while maintaining low queue delay.
@@ III-D Computation Overhead(s)
Figure 17 demonstrates LoRA’s parameter reduction efficiency: Llama2’s 7B parameters reduce to 70M (1%), while OPT’s 1.3B decrease to 25M (1.9%) using rank-128 LoRA. GPT-2 (125M) and T5-LLM (220M) need no reduction due to their manageable size. This rank-128 configuration optimally balances efficiency with performance, maintaining full fine-tuning accuracy while reducing computational demands. Figure 18 highlights memory constraints critical for deployment: Llama2 requires 23.91GB VRAM, OPT 7GB, GPT2 3.6GB, and T5-LLM just 2.32GB. These requirements emphasize the need for efficient architectures in resource-constrained networking environments where routers must operate within strict hardware limitations while maintaining real-time performance.
@@
Figure 19 demonstrates L4S-LLM’s significant impact on inference latency. The unadapted Llama2-7B exhibits prohibitive answer generation time of 0.200 seconds, while LoRA-adapted models achieve substantial reductions: Llama2 at 0.0391s, OPT at 0.0395s, GPT2 at 0.0414s, and T5-LLM at 0.0411s. Despite Llama2-7B and OPT-1.3B containing billions of parameters, LoRA reduces inference time by approximately 80%, achieving  0.039s response times crucial for real-time congestion control. This 5x speedup validates L4S-LLM’s efficiency for latency-critical network applications. However, further optimization opportunities exist through small language models (SLMs), model pruning, partitioning strategies, and specialized hardware acceleration to achieve sub-millisecond inference suitable for high-frequency packet processing.
@@
In Figure 20, we observe the total training time required to run 80 epochs for four different LLMs. Even with the application of LoRA, Llama2-7B exhibits the highest training time by a significant margin at approximately 22.3 hours. In contrast, the training time for OPT with LoRA is around 29.9 minutes, while GPT2 requires approximately 17.96 minutes, and T5-LLM requires about 32.26 minutes. With the exception of Llama2, the training times for other LLM models are notably faster. However, during evaluation or actual deployment, the inference time is of utmost importance. Despite Llama2’s higher training time, its inference performance is comparable to other models, making it suitable for real-time congestion control applications.
III-E Model Architecture and Optimization Considerations
Our architectural decisions were strategically formulated to achieve an optimal balance between performance requirements and deployment feasibility. The L4S-LLM framework implements a robust three-component architecture: (1) a sophisticated state encoder utilizing linear layers for scalar inputs and CNNs for time-series data, (2) a powerful transformer-based reasoning core, and (3) an innovative L4S-LLM head enabling single-inference action generation that significantly reduces decision latency.
Through extensive network evaluation, we systematically tested different foundation models (Llama2-7B, OPT-1.3B, GPT2-125M, T5-220M). While Llama2-7B demonstrated superior accuracy (97.56%), smaller models achieved a substantial 83% accuracy, providing compelling evidence for viable architectural simplification pathways that maintain acceptable performance levels.
Regarding the critical latency-complexity tradeoff, our measured decision times (39-41ms) currently exceed the ideal 15ms target for high-frequency AQM operations. We effectively mitigated this limitation through periodic execution at 10 and 100-packet intervals (Figures 11-16), conclusively demonstrating that even with reduced execution frequency, L4S-LLM delivers consistently superior performance compared to traditional approaches across all key metrics.
Our current implementation focuses on parameter-efficient fine-tuning through LoRA, achieving a remarkable 99% reduction in trainable parameters. We fully recognize the importance of additional optimization techniques for production deployment. Future work will systematically explore structured pruning to eliminate non-essential network components, quantization-aware training for lower precision operation, and hardware-aware optimization specifically targeting router-specific constraints. Based on recent advances in model compression techniques [10, 13], these approaches could potentially reduce model size by an additional 75-90% with minimal performance degradation, making deployment on router-grade hardware an achievable near-term goal.
Our comprehensive open-source FreeBSD-14 implementation establishes a solid foundation for the networking community to collaboratively explore these optimization approaches as we advance toward production-ready systems. Several critical limitations must be addressed before considering the commercial deployment of the proposed L4S-LLM framework in network routers. This section outlines the primary challenges affecting deployability and overall effectiveness:
  • Online Training/Testing: The framework requires real-time evaluation within the L4S architecture to address deployment challenges. While the current L4S-LLM framework was trained and tested using a data-driven approach with experimental data from L4S DualPI2 AQM and various congestion control algorithms, this process involved offline training and testing.
@@ III-E Model Architecture and Optimization Considerations
  • Energy: Continuous operation of LLMs requires substantial energy, typically supported by high-performance GPUs and CPUs. The significant energy demands and associated costs pose a challenge to commercial viability. Major technology firms are exploring nuclear-powered energy solutions to meet their LLM needs.
@@ IV Conclusion
This research introduces AQM-LLM, a framework that adapts large language models for active queue management within the L4S architecture. Leveraging preemptive ECN signalling and packet dropping, L4S-LLM dynamically controls congestion, optimizes queues, and minimizes latency. Low-rank adaptation cuts memory use by 64%, while a data-driven reinforcement learning pipeline streamlines fine-tuning. An open-source L4S-LLM platform on FreeBSD-14 supports future research and IETF standardization. While delivering major gains in latency, throughput, and efficiency, further refinement via pruning, quantization, and distillation toward small-language-model-based AQM can enhance scalability and resource efficiency.
Appendix
Towards L4S-LLM Stability and Convergence
We analyze the stability and convergence properties of L4S-LLM by integrating transformer-based language model theory with control-theoretic principles of L4S AQM systems. Considering transformer dynamics and convergence, the self-attention mechanism in L4S-LLM can be approximated as
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V.$ (5)
where $Q,K,V\in\mathbb{R}^{n\times d_{k}}$ represent query, key, and value matrices derived from network state embeddings. For network state sequences $s=(s_{1},s_{2},\ldots,s_{n})$, the attention mechanism creates a weighted representation that captures temporal dependencies across network metrics. The transformer layer mapping $F$ exhibits convergence when its Lipschitz constant satisfies
$\mathcal{L}=\sup_{h\neq h^{\prime}}\frac{\|F(h)-F(h^{\prime})\|}{\|h-h^{\prime}\|}<1.$ (6)
With LoRA adaptation, the effective weight matrices become $W=W_{0}+AB$ where $W_{0}$ is the pre-trained weight, and $A\in\mathbb{R}^{d\times r}$, $B\in\mathbb{R}^{r\times k}$ are low-rank adaptation matrices. The stability of this adapted system depends on the spectral radius of the Jacobian:
$\rho\left(\frac{\partial F_{\text{LoRA}}(h)}{\partial h}\right)<1$ (7)
where $F_{\text{LoRA}}$ is the LoRA-adapted transformer mapping. Our empirical results with rank $r=128$ confirm this stability condition is maintained.
Employing L4S control theoretic approach, the L4S-LLM policy $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}$ replaces the traditional PI controller dynamics
$p^{\prime}(t)=p^{\prime}(t-1)+\alpha(q_{\text{delay}}(t)-q_{\text{target}})+\beta(q_{\text{delay}}(t)-q_{\text{delay}}(t-1)).$
Given Lyapunov function $V(s)=(q_{\text{delay}}-q_{\text{target}})^{2}$, stability is guaranteed when
$\mathbb{E}[V(s_{t+1})|s_{t},\pi_{\theta}(s_{t})]-V(s_{t})<0,\quad\forall s_{t}\neq s^{*}$ (8)
For the RL-optimized policy, convergence follows
$\|J(\pi_{k})-J(\pi^{*})\|\leq(1-\gamma\lambda)^{k}\|J(\pi_{0})-J(\pi^{*})\|$ (9)
where $J(\pi)$ is the expected return, $\pi^{*}$ is the optimal policy, and $\lambda$ relates to the mixing time of the induced Markov chain. Therefore, L4S-LLM provides probabilistic queue stability
$\mathbb{P}\left(\sup_{t\geq T_{0}}|q_{\text{delay}}(t)-q_{\text{target}}|\leq\epsilon\right)\geq 1-\delta$ (10)
for finite $T_{0}$, stability bound $\epsilon$, and error probability $\delta$.
Observe that this theoretical analysis is substantiated by the experimental results discussed earlier, for examples, rapid convergence of training loss (Figure 9), high accuracy in action prediction (97.56% for Llama2), significant improvements in queue delay stability (Figure 11), enhanced bandwidth utilization (Figure 14). Our understanding is that the integration of LLM and ha the potential to yield an enhanced AQM system with strong stability and convergence properties, capable of outperforming traditional approaches.
Model Compressions for Router-Deployable L4S-LLM
To address L4S-LLM deployment’s computational limitations, we propose four optimization strategies to reduce model size while maintaining performance:
Structured Pruning
Using attribution scores and confidence-weighted patterns, approximately 40% of attention heads could be eliminated with negligible impact, as network traffic patterns exhibit strong temporal locality. Selective pruning of FFNs (65% of parameters) based on activation patterns during congestion events could reduce dimensions by 50-60% while maintaining responsiveness. Layer-specific compression rates guided by sensitivity analysis would preserve parameter density in layers with greater impact on queue management.
Quantization-Aware Training
Critical components would maintain FP16 precision, while less sensitive components could operate at INT8/INT4, reducing memory requirements by up to 75%. Quantization parameters would be calibrated using diverse traffic patterns, emphasizing accuracy during congestion transition states. Adaptive precision would allocate higher bit-width to components processing critical congestion indicators while using lower precision for stable states.
Neural Architecture Search
Evolutionary search algorithms would explore architectural variants optimized for congestion control. Asymmetric encoder designs with wider attention layers but narrower feedforward networks could maintain 95% performance while reducing parameters by 70%. Custom components for network telemetry could replace general-purpose transformers, with dedicated convolutional pathways for multi-timescale queue dynamics.
Hardware-Aware Optimization
Critical inference paths would be optimized to reduce memory transfers for efficient execution on processors with limited cache. The model would develop structured sparsity patterns aligned with hardware execution blocks. Router-specific optimizations would include operator reordering, buffer reuse, and execution scheduling tailored to hardware constraints.